OpenStudy (anonymous):

A land developer sold 4 1/2 acres of his 15 1/5 acres. How much does he have left?

OpenStudy (calculusxy):

\[15 \frac{1}{5} -(15 \frac{ 1 }{ 5 } \times 4\frac{ 1 }{ 2 })\]